Construct a 99​%confidence interval of the population proportion using the given information.

x equals 160 comma n equals 200x=160, n=200
Answer #1

Solution :

Given that,

n = 200

x = 160

Point estimate = sample proportion = = x / n = 160/200=0.8

1 -   = 1- 0.8 =0.2

At 99% confidence level the z is ,

  = 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.576 ( Using z table )

  Margin of error = E = Z/2   * (((( * (1 - )) / n)

= 2.576* (((0.8*0.2) /200 )

= 0.0729

A 99% confidence interval for population proportion p is ,

- E < p < + E

0.8-0.0729 < p <0.8+ 0.0729

0.7271< p < 0.8729

The 99% confidence interval for the population proportion p is : 0.7271, 0.8729